# Quantum many-body scars with unconventional superconducting pairing symmetries via multi-body interactions

### v2

@@ VI.5 Higher multi-body interactions
Figure 11: Energy cancellation between the two-body on-site interaction (red box) and the multi-body interaction acting among a doublon and two neighboring particles (blue box).
@@
To achieve superconducting scar states, it is also possible to exploit higher multi-body interactions (i.e., $M$-body interactions with larger $M$). In the case of the $d$-wave $\eta$-pairing state on the square lattice (discussed in Sec. III), we find that a doublon created by an overlap between two pairs is always accompanied by two particles around the doublon. We can cancel the on-site two-body interaction with a multi-body interaction acting among a doublon and two associated particles, as shown by the blue box in Fig. 11. Motivated by this observation, we consider the following interaction term in the Hamiltonian,
@@ VI.6 Pair number operator
Using the neighboring particle existence operator $\mathbb{n}_{i\sigma}$ [Eq. (25)], we can measure the number of pairs correctly, as shown in Fig. 12(b). The $\eta$-pairing states become an eigenstate of $\hat{N}_{p}$:
@@
The operator $\hat{N}_{\mathrm{p}}$ is used to stabilize the unconventional $\eta$-pairing states over unpaired states. Figure 12(c) shows an example of a Fock state with one broken pair. The expectation value of $\hat{N}_{\mathrm{p}}$ for this state is $3/2$. If we introduce a pair chemical potential $\mu_{\mathrm{p}}$ and replace the Hamiltonian $\mathcal{H}_{\mathrm{extH}}$ with $\mathcal{H}_{\mathrm{extH}}-\mu_{\mathrm{p}}\hat{N}_{\mathrm{p}}$, we can decrease the eigenenergy of the unconventional pairing states than other unpaired eigenstates.
@@ VI.7 Open boundary condition
The interaction term $\mathcal{H}_{\rm int}$ of the Hamiltonian in the open boundary condition is defined as in Eq. (30) with the neighboring particle existence operator $\mathbb{n}_{i\sigma}$ given by Eq. (26), where the coordination number $z$ changes at the boundary. One can confirm that the off-site $\eta$-pairing states are eigenstates of the interaction term $\mathcal{H}_{\mathrm{int}}$ with the appropriately defined $\mathbb{n}_{i\sigma}$ in the open boundary condition. As a result, the spin-triplet $d$-wave $\eta$-pairing states given in Eq. (19), for instance, are the exact energy eigenstates of $\mathcal{H}_{\mathrm{extH}}$ [Eq. (29)] even in the open boundary condition.
@@ VII Summary and outlook
In this paper, we have presented a systematic framework to construct model Hamiltonians that have superconducting quantum many-body scar states with unconventional pairing symmetries, e.g., the $s$-wave spin-triplet (or spinless), $p$-wave spin-singlet, $d$-wave spin-triplet, and $f$-wave spin-triplet symmetries, by engineering multi-body interactions. The key idea was to cancel the energy increment due to the formation of doublons by multi-body interactions. To this end, we have introduced the neighboring particle existence operator $\mathbb{n}_{i\sigma}$ in Eq. (25), which judges whether there exist spin-$\sigma$ particles on the nearest-neighbor sites of site $i$. Using this operator, we have defined the multi-body interaction terms in the form of $\sum_{\sigma}n_{i\uparrow}n_{i\downarrow}\mathbb{n}_{i\sigma}$ in Eq. (28), which precisely cancel the two-body interactions acting among off-site pairs (while working non-trivially on unpaired particles). We have applied our approach to the two-dimensional extended Hubbard model with the multi-body interactions, and numerically confirmed that the spin-triplet $d$-wave $\eta$-pairing state is indeed the QMBS eigenstate. Our construction is flexible enough that it can be applied to various pairing symmetries and arbitrary bipartite lattice models, some of which are discussed in the paper (see Table 1). The derived unconventional pairing states are natural extensions of Yang’s $s$-wave spin-single $\eta$-pairing states, and will be relevant for applications to nonthermal and long-lived nonequilibrium superconductivity with unconventional pairing symmetries.
@@
There are various future problems. From a practical point of view, it will be important to find a protocol to reach the unconventional superconducting QMBS states from experimentally accessible initial states. For the conventional $s$-wave $\eta$-pairing states, there have been several proposals including the periodic drive [79, 80, 81, 82], photo-doping [83, 19, 84, 85, 86, 87, 68], dissipation engineering [88, 89, 90, 91], and adiabatic driving [92], some of which could be extended to the unconventional cases. Another issue is the stability of the unconventional superconducting scar states against various perturbations [59, 93]. In our models, we need to fine tune the interaction parameters. It is necessary to understand how the scar states can survive (or not) when the values of those parameters deviate from the optimal ones. Topological aspects of the scar states (such as the $p_{x}+ip_{y}$-wave $\eta$-pairing state discussed in Sec. VI.1) will be another interesting topic, where the question is whether a bulk-boundary correspondence also holds for quantum many-body scar states. It will also be interesting to explore the relation between the present unconventional superconducting scar states and the fracton physics [94, 95, 96, 35, 64], since our model exhibits a high degree of degeneracy of the eigenstates which grow as a function of the system size.
@@
Finally, we briefly comment on possible experimental realizations of the superconducting QMBS states with unconventional pairing symmetries. It is required that the scarred models have a controllable multi-body interaction. The spinless/spin-polarized $s$-wave pairing states discussed in Secs. III.1 and VI.4 can be implemented in the one-dimensional system with the density-density-type three-body interaction in Eq. (12). The three-body interaction has been realized in a controllable manner in cold polar-molecular gases trapped in an optical lattice by means of effective interactions mediated by dipole-dipole interactions [97, 98, 99, 100, 101, 102]. Thus, the unconventional superconducting scar state in Eq. (9) in the system described by Eq. (13) is a primary example of what can be achieved with the state-of-the-art experimental techniques for ultracold atomic systems.

### v3

@@ VI.5 Higher multibody interactions
Figure 11: Energy cancellation between the two-body on-site interaction (red box) and the multibody interaction acting among a doublon and two neighboring particles (blue box).
@@
To achieve superconducting scar states, it is also possible to exploit higher multibody interactions (i.e., $M$-body interactions with larger $M$). In the case of the $d$-wave $\eta$-pairing state on the square lattice (discussed in Sec. III), we find that a doublon created by an overlap between two pairs is always accompanied by two particles around the doublon. We can cancel the on-site two-body interaction with a multibody interaction acting among a doublon and two associated particles, as shown by the blue box in Fig. 11. Motivated by this observation, we consider the following interaction term in the Hamiltonian,
@@ VI.6 Pair number operator
Using the neighboring particle existence operator $\mathbb{n}_{i\sigma}$ [Eq. (25)], we can measure the number of pairs correctly, as shown in Fig. 12(b). The $\eta$-pairing states become an eigenstate of $\hat{N}_{\mathrm{p}}$:
@@
The operator $\hat{N}_{\mathrm{p}}$ is used to stabilize the unconventional $\eta$-pairing states over unpaired states. Figure 12(c) shows an example of a Fock state with one broken pair. The expectation value of $\hat{N}_{\mathrm{p}}$ for this state is $3/2$. If we introduce a pair chemical potential $\mu_{\mathrm{p}}$ and replace the Hamiltonian $\mathcal{H}_{\mathrm{extH}}$ with $\mathcal{H}_{\mathrm{extH}}-\mu_{\mathrm{p}}\hat{N}_{\mathrm{p}}$, we can decrease the eigenenergy of the unconventional pairing states than other unpaired eigenstates. Since the pair chemical potential $\mu_{\mathrm{p}}$ lowers the energy of the entire subspace where all particles form pairs (not limited to the unconventional $\eta$-pairing states), an important remaining question is whether superconductivity may emerge in the ground state within this subspace.
@@ VI.7 Open boundary condition
The interaction term $\mathcal{H}_{\mathrm{int}}$ of the Hamiltonian in the open boundary condition is defined as in Eq. (30) with the neighboring particle existence operator $\mathbb{n}_{i\sigma}$ given by Eq. (26), where the coordination number $z$ changes at the boundary. One can confirm that the off-site $\eta$-pairing states are eigenstates of the interaction term $\mathcal{H}_{\mathrm{int}}$ with the appropriately defined $\mathbb{n}_{i\sigma}$ in the open boundary condition. As a result, the spin-triplet $d$-wave $\eta$-pairing states given in Eq. (19), for instance, are the exact energy eigenstates of $\mathcal{H}_{\mathrm{extH}}$ [Eq. (29)] even in the open boundary condition.
@@ VII Summary and outlook
In this paper, we have presented a systematic framework to construct model Hamiltonians that have superconducting quantum many-body scar states with unconventional pairing symmetries, e.g., the $s$-wave spin-triplet (or spinless), $p$-wave spin-singlet, $d$-wave spin-triplet, and $f$-wave spin-triplet symmetries, by engineering multibody interactions. The key idea was to cancel the energy increment due to the formation of doublons by multibody interactions. To this end, we have introduced the neighboring particle existence operator $\mathbb{n}_{i\sigma}$ in Eq. (25), which judges whether there exist spin-$\sigma$ particles on the nearest-neighbor sites of site $i$. Using this operator, we have defined the multibody interaction terms in the form of $\sum_{\sigma}n_{i\uparrow}n_{i\downarrow}\mathbb{n}_{i\sigma}$ in Eq. (28), which precisely cancel the two-body interactions acting among off-site pairs (while working non-trivially on unpaired particles). We have applied our approach to the two-dimensional extended Hubbard model with the multibody interactions, and numerically confirmed that the spin-triplet $d$-wave $\eta$-pairing state is indeed the QMBS eigenstate. Our construction is flexible enough that it can be applied to various pairing symmetries and arbitrary bipartite lattice models, some of which are discussed in the paper (see Table 1). The derived unconventional pairing states are natural extensions of Yang’s $s$-wave spin-single $\eta$-pairing states, and will be relevant for applications to nonthermal and long-lived nonequilibrium superconductivity with unconventional pairing symmetries that might be difficult to achieve in thermal equilibrium.
@@
There are various future problems. From a practical point of view, it will be important to find a protocol to reach the unconventional superconducting QMBS states from experimentally accessible initial states. For the conventional $s$-wave $\eta$-pairing states, there have been several proposals including the periodic drive [87, 88, 89, 90], photo-doping [91, 19, 92, 93, 94, 95, 68], dissipation engineering [96, 97, 98, 99], and adiabatic driving [100], some of which could be extended to the unconventional cases. Another issue is the stability of the unconventional superconducting scar states against various perturbations [59, 101]. It is necessary to understand how the scar states can persist (or not) when the parameters of the multibody interactions deviate from the ones we obtained. Furthermore, it is known that Yang’s $\eta$-pairing state exhibits electromagnetic instability in the presence of external electromagnetic fields [60, 61]. However, in the case of unconventional superconducting scar states, $\eta$-pairing states with different (odd) parity can potentially lead to qualitatively distinct electromagnetic stability, which warrants further investigation. Topological aspects of the scar states (such as the $p_{x}+ip_{y}$-wave $\eta$-pairing state discussed in Sec. VI.1) will be another interesting topic, where the question is whether a bulk-boundary correspondence also holds for quantum many-body scar states. It will also be interesting to explore the relation between the present unconventional superconducting scar states and the fracton physics [102, 103, 104, 35, 64], since our model exhibits a high degree of degeneracy of the eigenstates which grow as a function of the system size.
@@
Finally, we briefly comment on possible experimental realizations of the superconducting QMBS states with unconventional pairing symmetries. It is required that the scarred models have a controllable multibody interaction. The spinless/spin-polarized $s$-wave pairing states discussed in Secs. III.1 and VI.4 can be implemented in the one-dimensional system with the density-density-type three-body interaction in Eq. (12). The three-body interaction has been realized in a controllable manner in cold polar-molecular gases trapped in an optical lattice by means of effective interactions mediated by dipole-dipole interactions [105, 106, 107, 108, 109, 110]. Thus, the unconventional superconducting scar state in Eq. (9) in the system described by Eq. (13) is a primary example of what can be achieved with the state-of-the-art experimental techniques for ultracold atomic systems.